A stone is thrown straight downward with a speed of 20 m/s from the top of a tall building. If the stone strikes the ground 3 s later, how tall is the building? Assume air resistance is negligible.


1
Expert's answer
2022-01-23T16:14:45-0500

Given:

"v_0=20\:\rm m/s"

"g=9.8\:\rm m/s^2"

"t=3.0\:\rm s"


The height of a building

"h=v_0t+gt^2/2\\
=20*3.0+9.8*3.0^2/2=104\:\rm m"